Star Cups, a specialty coffee store, creates a special mix of two coffees at two different prices $6.42 and $7.28 per pound each. How many pounds of the $7.28 coffee should be mixed with 7 pounds of $6.42 coffee to sell the mixture for $7.20 per pound?
x+7 pounds = total quanity of mixture.
6.42(7) + 7.28x = 7.20 (x+7)
642(7)+728=720 (x+7)
4494 + 728x = 720x +5040
4494 +728x=720x + 5040
4494 +728x-720x=720x-720=5040
28x =546
7x/7x=546/7
x=72

Answer by checkley75(3666)

7*6.42+7.28x=7.20(x+7)
44.94+7.28x=7.20x+50.40
7.28x-7.20x=50.40-44.94
.08x=5.46
x=5.46/.08
x=68.25 pounds of $7.28 coffee is needed
proof
7*6.42+7.28*68.25=7.20*75.25
44.94+496.86=541.8
541.8=541.8
----------------------------------------
you have too many = signas & you drop the x factor in your calculations.
you need to be more careful in your tranferring your values in each step.